OpenStudy (emilyjones284):

Please help...Ill give a medal (: Ill post the question below (:

OpenStudy (anonymous):

TELL ME THE QUESTION!! >:D

OpenStudy (emilyjones284):

Solve \[8+5^{x}=1008\] Round to the nearest ten-thousandth.

OpenStudy (anonymous):

I got this.. x=200

OpenStudy (emilyjones284):

haha darn i would've given you a medal :(

OpenStudy (anonymous):

what @escrite said

OpenStudy (anonymous):

shes right

OpenStudy (anonymous):

give her a medal for being right

OpenStudy (emilyjones284):

would you mind showing me how you got that

OpenStudy (anonymous):

8 + 5x = 1008 8 + -8 + 5x = 1008 + -8 8 + -8 = 0 0 + 5x = 1008 + -8 5x = 1008 + -8 1008 + -8 = 1000 5x = 1000 Divide each side by '5'. x = 200

OpenStudy (anonymous):

amazingness is unraveling before my eyes

OpenStudy (emilyjones284):

Okay i probably should've mentioned that we"re using logs right now so are you sure thats right?

OpenStudy (emilyjones284):

and its not 5x ...the exponent is x

OpenStudy (anonymous):

Ah. That complicates the problem. You should have written x as an exponent then. 8+5^x=1008

OpenStudy (emilyjones284):

haha i did write it as an exponent

OpenStudy (anonymous):

lol she did

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

I see that now, my poor eyesight is making me look dumb cx

OpenStudy (emilyjones284):

Haha no its ok (: do you know how to solve it?

OpenStudy (anonymous):

I can give it a shot x)

OpenStudy (emilyjones284):

Okay thanks that would be greatly appreciated haha

OpenStudy (emilyjones284):

Hey can someone pleaseeeeeee help me haha (:

OpenStudy (anonymous):

4.29202967422018, got this from using the logarithmic function of a calculator

OpenStudy (anonymous):